40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  CTN Lab elog  Not logged in ELOG logo
Entry  Wed Feb 7 21:47:44 2018, Craig, awade, DailyProgress, NoiseBudget, February 7th Noisebudget 20180207_221703noiseBudget.pdf
    Reply  Thu Feb 8 18:38:04 2018, rana, DailyProgress, NoiseBudget, February 7th Noisebudget 
       Reply  Fri Feb 9 11:36:08 2018, awade, DailyProgress, NoiseBudget, 200-3000 Hz hump search 
          Reply  Mon Feb 12 16:29:50 2018, awade, Craig, DailyProgress, NoiseBudget, 200-3000 Hz hump search 
             Reply  Fri Feb 16 12:41:32 2018, awade, Craig, DailyProgress, NoiseBudget, 200-3000 Hz hump search Beatnote_ASD_20180215_121922.pdfBeatnote_ASD_20180215_193422.pdfBeatnote_CurrentASD.pdf
Message ID: 2087     Entry time: Mon Feb 12 16:29:50 2018     In reply to: 2083     Reply to this: 2094
Author: awade, Craig 
Type: DailyProgress 
Category: NoiseBudget 
Subject: 200-3000 Hz hump search 

Attempting to identify the source of the 200-3000kHz hump.  

Optimizing polarization into north BBEOM and AEOM

I made a number of optimizations of polarization into the modulating optics over the weekend.  The BBEOM used for fast actuation of the north FSS loop was optimized by driving the BBEOM with a 20 Vpp @ 101 Hz sin wave and minimizing the residual AM (as seen on the PMC reflection PD).

For the amplitude modulator used for the ISS RIN suppression, I deactivated this optically.  The polarization was aligned horizontal going into the AEOM and the following quarter-wave plate was set for maximum throughput.  In this configuration applied voltage to the AEOM results in little to no actuation.  I did this to remove the AEOM from consideration in diagnosing the 14.75 MHz RFAM residual.  

Attempting to improve the north AM residual from the 14.75 MHz modulation

I found that the residual AM, as see by the FSS Relf PD, was considerably higher than it could or should be. The initial level of AM was about -30 dBm.  With some adjustment of the two wave plates before the EOM I found I could reduce this level to -37 dBm. However, this was still a long way from the previous <-70 dBm achieved in the past.

After checking I was actually using a lambda/2 + lambda/4 pair I found that the pointing through the EOM (NP 4004) itself had an impact on the degree of AM reduction. It turns out that the strait shoot through the center of the aperture had a higher minimum RFAM residual compared to aligning the beam a little to the left on the output aperture (looking along the axis of the device). The beam is still well clear of the aperture but I was able to get to -72 dBm of residual AM.

What I found was that the AM would reduce for a minute or two and then creep back up to ~-40 dBm.  I suspected that this was temperature related. I pointed a heat gun at the EOM+ wave plates area at a distance of 1.5 meter for 5 seconds: it blew the AM reduction back up to -30 dBm.  It didn't seem to matter if the thermal hat was on the EOM or not.  I'm pretty sure these wave plates are zero-order but not sure how to tell just by looking at the optic, its possible that the label on the mount doesn't tell the truth.  

When I tried this optimization again today I had a lot of trouble finding an and optimal input polarization.  Eventually I've identified a source of polarization drift that is occurring somewhere between the 21.5 MHz PMC modulator and just before the wave plates into the 14.75 MHz oscillator.  This path is basically a two steering mirrors and the PMC.  I put a PBS in transmission between the PMC and the input to the north 14.75 MHz EOM. I found that the stability of the polarization alignment into that EOM improved significantly.  It was possible to get the residual AM to dip down to the -80 dBm noise floor of the analyzer (for that BW setting). There is, however, still a strong thermal dependence somewhere there which leads to the residual RFAM rising again.  

Side note: the orders for PMC mirrors indicate that p-pol mirrors were used.  Evan's notes don't seem justify this choice. I am working on the assumption that the PMCs should be operated using p- polarized light.

 

 

 

ELOG V3.1.3-